Solve each system by substitution. Determine whether the equations are independent, dependent or inconsistent:
y= -3x+19
y= 2x-1

Answer by stanbon(57203)

Solve each system by substitution. Determine whether the equations are independent, dependent or inconsistent:
y= -3x+19
y= 2x-1
-------------
2x-1 = -3x+19
5x = 20
x = 4
------
If x=4, then y = 2*4-1 = 7
----------
System solution: (4,7)
Therefore the system is consistent.
=======================================
Cheers,
Stan H.